OpenStudy (anonymous):

Could someone so kindly help me with a problem? I guess I missed this part of the lesson I have no idea what this is. I need to simplify this: e ^{3ln x}

OpenStudy (anonymous):

\[e ^{3\ln x}\]

ganeshie8 (ganeshie8):

you need to use below properties : \(a \log b = \log b^a\)

ganeshie8 (ganeshie8):

and one more : \(\large e^{\ln x} = x\)

ganeshie8 (ganeshie8):

maybe u can ponder why they work later, but for now u can simplify ur expression using these

OpenStudy (whpalmer4):

\[e^{3\ln x} = e^{\ln x}*e^{\ln x}*e^{\ln x}\]

OpenStudy (anonymous):

awesome this helps so much thank you guys :)

ganeshie8 (ganeshie8):

np :)